### v1

# Asynchronous Federated Reinforcement Learning with Policy Gradient Updates: Algorithm Design and Convergence Analysis

Abstract.
To improve the efficiency of reinforcement learning, we propose a novel asynchronous federated reinforcement learning framework termed AFedPG, which constructs a global model through collaboration among $N$ agents using policy gradient (PG) updates. To handle the challenge of lagged policies in asynchronous settings, we design delay-adaptive lookahead and normalized update techniques that can effectively handle the heterogeneous arrival times of policy gradients. We analyze the theoretical global convergence bound of AFedPG, and characterize the advantage of the proposed algorithm in terms of both the sample complexity and time complexity. Specifically, our AFedPG method achieves $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ sample complexity at each agent on average. Compared to the single agent setting with $\mathcal{O}(\epsilon^{-2.5})$ sample complexity, it enjoys a linear speedup with respect to the number of agents. Moreover, compared to synchronous FedPG, AFedPG improves the time complexity from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$, where $t_{i}$ denotes the time consumption in each iteration at the agent $i$, and $t_{\max}$ is the largest one. The latter complexity $\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$ is always smaller than the former one, and this improvement becomes significant in large-scale federated settings with heterogeneous computing powers ($t_{\max}\gg t_{\min}$). Finally, we empirically verify the improved performances of AFedPG in three MuJoCo environments with varying numbers of agents. We also demonstrate the improvements with different computing heterogeneity.
Reinforcement Learning, Policy Gradient, Federated Learning, Asynchronous, Optimization
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; 2024; isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence Distributed artificial intelligenceccs: Computer systems organization Distributed architecturesccs: Networks Network services
1. Introduction
Policy gradient (PG) methods, also known as REINFORCE (Williams, 1992), are widely used to solve reinforcement learning (RL) problems across a range of applications, with recent notable examples including reinforcement learning from human feedback (RLHF) in Google’s Gemma (Team and DeepMind, 2024), OpenAI’s InstructGPT (Ouyang et al., 2022), and ChatGPT (GPT-4 (OpenAI, 2023)). It has also garnered significant attention in communications/networking communities for solving network optimization (Wei et al., 2023; Hu et al., 2022) and wireless communication (Yu et al., 2023; Xiong et al., 2022) problems.
Most practical RL applications are on large scales and rely on a huge amount of data samples for model training in an online behavior (Provost and Fawcett, 2013; Liu et al., 2021), which is considered as the key bottleneck of RL (Dulac-Arnold et al., 2021; Ladosz et al., 2022). To reduce sample complexity while enabling training on massive data, a conventional approach involves transmitting locally collected samples from distributed agents to a central server (Predd et al., 2006). The transmitted samples can then be used for policy learning on the server side. However, this may not be feasible in real-world mobile systems where the communication bandwidth is limited and long delays are intolerable (Hu et al., 2024; Lan et al., 2023a; Niknam et al., 2020), such as wireless edge devices (Han et al., 2023; Lan et al., 2023a), Internet of Things (Nguyen et al., 2021; Fang et al., 2023), autonomous driving (Kiran et al., 2022), and vehicle transportation (Al-Abbasi et al., 2019; Haliem et al., 2021). In RL, since new data is continuously generated based on the current policy, the samples collected at the agents need to be transmitted frequently to the server throughout the training process, resulting in a significant bottleneck with long delays. Furthermore, the sharing of individual data collected by agents may lead to privacy and legal issues (Kairouz et al., 2021; Mothukuri et al., 2021).
Federated learning (FL) (Hosseinalipour et al., 2020; McMahan et al., 2017) offers a promising solution to the above challenges in a distributed setting. In FL, instead of directly transmitting the raw datasets to the central server, agents only communicate locally trained model parameters (or gradients) with the server. Although FL has been mostly studied for supervised learning problems, several recent works expanded the scope of FL to federated reinforcement learning (FedRL), where $N$ agents collaboratively learn a global policy without sharing the trajectories they collected during agent-environment interaction (Jin et al., 2022; Khodadadian et al., 2022; Xie and Song, 2023). FedRL has been studied in the tabular case (Agarwal et al., 2021a; Jin et al., 2022; Khodadadian et al., 2022), and for value-function based algorithms (Wang et al., 2023; Xie and Song, 2023), where a linear speedup has been demonstrated. For policy-gradient based algorithms, namely FedPG, we note that linear speedup is easy to see as the trajectories collected by each agent could be parallelized (Lan et al., 2023b). However, policy-gradient based FedRL has not been well studied in terms of global (non-local) convergence and time complexity.
Despite all the aforementioned works in FedRL, they still face challenges in terms of time complexity, primarily due to their focus on synchronous model aggregation. In large-scale heterogeneous settings (Xie et al., 2019; Chen et al., 2020), performing synchronous global updates has limitations, as the overall time consumption heavily depends on the slow agents, i.e., stragglers (Mishchenko et al., 2022). In this paper, we aim to tackle this issue by strategically leveraging asynchronous federated learning (A-FL) in policy-based FedRL for the first time. A-FL (Xie et al., 2019) shows superiority compared to synchronous FL, and recent works (Mishchenko et al., 2022; Koloskova et al., 2022) further improve convergence performances with theoretical guarantees. However, compared to prior A-FL approaches focusing on supervised learning, integrating A-FL with policy-based FedRL introduces new challenges due to the presence of lagged policies in asynchronous settings. Unlike supervised FL where the datasets of the clients are fixed, in RL, agents collect new samples in each iteration based on the current policy. This dynamic nature of the data collection process makes both the problem itself and the theoretical analysis challenging. To the best of our knowledge, this problem setting and its challenges have been largely overlooked in existing research, despite the significance of employing FL in RL. The key question that this paper aims to address is:
Table 1. Performance improvements of our AFedPG over other policy gradient methods. We compare the complexity for global convergence in each agent.
PG FedPG AFedPG (ours)
Sample Complexity $\mathcal{O}({\epsilon}^{-2.5})$ $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$
Global Time - $\mathcal{O}(\frac{t_{\max}}{N}{\epsilon}^{-2.5})$ $\mathcal{O}(\bar{t}{\epsilon}^{-2.5})$
Despite the inherent challenge of dealing with lagged policies among different agents, can we improve the efficiency of FedPG through asynchronous methods while ensuring theoretical convergence?
We answer this question in the affirmative by proposing AFedPG, an algorithm that asynchronously updates the global policy using policy gradients from federated agents. The key components of the proposed AFedPG include a delay-adaptive lookahead technique, which addresses the inconsistent arrival times of updates during the training process. The improvement of AFedPG over other approaches is summarized in Table 1. Here, the global time is measured by the number of global updates $\times$ time complexity in each iteration. In terms of sample complexity, the federated learning technique brings a linear speedup with respect to the number of agents $N$. As for the global time, AFedPG improves from $\mathcal{O}(\frac{t_{\max}}{N}{\epsilon}^{-2.5})$ to $\mathcal{O}(\bar{t}{\epsilon}^{-2.5})$, where $\bar{t}\coloneqq 1/\sum_{i=1}^{N}\frac{1}{t_{i}}$ is a harmonic average and always smaller than $t_{\max}/N$, and $t_{i}$ denotes the time complexity in each iteration at agent $i$.
1.1. Summary of Contributions
Our main contributions can be summarized as follows:
  1. (1)
    New methodology with a delay-adaptive technique: We propose AFedPG, an asynchronous training method tailored to FedRL. To handle the delay issue in the asynchronous FedRL setting, we design a delay-adaptive lookahead technique. Specifically, in the $k$-th iteration of training, the agent collects samples according to the local model parameters $\widetilde{\theta}_{k}\leftarrow\theta_{k}+\frac{1-\alpha_{k-\delta_{k}}}{\alpha_{k-\delta_{k}}}(\theta_{k}-\theta_{k-1})$, where $\delta_{k}$ is the delay. This updating technique cancels out the second-order correction terms and thus assists the convergence analysis.
  2. (2)
    Convergence analysis: This work gives the global convergence guarantees of the federated policy-based RL for the first time. We analytically characterize the convergence bound of AFedPG using the key lemmas, and show the impact of various parameters including delay and number of iterations in our analysis.
  3. (3)
    Linear speedup in sample complexity: As shown in Table 1, our federated method improves the sample complexity in each agent from $\mathcal{O}({\epsilon}^{-2.5})$ (single agent PG) to $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$, where $N$ is the number of federated agents. This represents the linear speedup with respect to the number of agents $N$.
  4. (4)
    Time complexity improvement: Our AFedPG also reduces the time complexity of synchronous FedPG from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}(\bar{t}\coloneqq\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$. The latter is always smaller than the former. This gap is significant in large-scale federated settings with heterogeneous computing powers ($t_{\max}\gg t_{\min}$).
  5. (5)
    Experiments under MuJoCo environment: We empirically verify the improved performances of AFedPG in three different MuJoCo environments with varying numbers of agents. We also demonstrate the improvements with different computing heterogeneity.
To the best of our knowledge, this is the first work to successfully integrate policy-based reinforcement learning with asynchronous federated learning and analyze its behavior. This new setting necessitates us to deal with the lagged policies under a time-varying data scenario depending on the current policy.
[…]
The remainder of this paper is organized as follows. In the next subsection, we summarize the related works. In Section 2, we introduce our problem setting, including some backgrounds on reinforcement learning and policy gradients. In Section 3, we describe our proposed AFedPG method. In Section 4, we analyze the global convergence rate of AFedPG and discuss its time complexity compared to the synchronous setting. Section 5 presents the experimental results, and we conclude this paper in Section 6.
1.2. Related Work
In this subsection, we review previous works that are most relevant to the present work.
Policy gradient methods: For vanilla PG, the state-of-the-art result is presented in (Yuan et al., 2022), achieving a sample complexity of $\widetilde{\mathcal{O}}({\epsilon}^{-4})$ for the local convergence. Several recent works have improved this boundary with PG variants. In (Huang et al., 2020), a PG with momentum method is proposed with convergence rate $\mathcal{O}({\epsilon}^{-3})$ for the local convergence. The authors of (Ding et al., 2022) further improve the convergence analysis of PG with momentum and achieve a global convergence with the rate $\mathcal{O}({\epsilon}^{-3})$. In (Fatkhullin et al., 2023), a normalized PG technique is introduced and improves the global convergence rate to $\mathcal{O}({\epsilon}^{-2.5})$. In (Mondal and Aggarwal, 2024), an acceleration-based natural policy gradient method is proposed with sample complexity $\mathcal{O}({\epsilon}^{-2})$, but second-order matrices are computed in each iteration (with first-order information), which brings more computational cost. However, all previous works have focused on the single-agent scenario, and the federated PG has not been explored. Compared to these works, we focus on a federated PG setting to improve the efficiency of RL in practical settings with distributed agents. Our scheme achieves linear speedup with respect to the number of agents, significantly reducing the sample complexity compared to the conventional PG approaches.
Asynchronous FL: In (Xie et al., 2019), the superiority of asynchronous federated learning (A-FL) has been empirically shown compared to synchronous FL in terms of convergence performances. In (Chen et al., 2020), the asynchronous analysis is extended to the vertical FL with a better convergence performance compared to the synchronous setting. In (Dun et al., 2023), a dropout regularization method is introduced to handle the heterogeneous problems in A-FL. About the same time, recent works (Mishchenko et al., 2022; Koloskova et al., 2022) further improve the convergence performance with theoretical guarantees, and show that A-FL always exceeds synchronous FL without any changes to the algorithm. We note that all previous A-FL works focus on supervised learning with fixed datasets on the client side. However, in RL, agents collect new samples that depend on the current policy (model parameters) in each iteration, which makes the problem fundamentally different and challenging. In this work, we address this challenge by developing an A-FL method highly tailored to policy gradient.
FedRL: For value-function based algorithms, (Jin et al., 2022) analyzes the convergence performances with environment heterogeneity. In (Agarwal et al., 2021a), a local update method is used to reduce communication costs. (Khodadadian et al., 2022) shows a linear speedup with respect to the number of agents. All of the above works are limited to tabular RL analysis. For actor-critic based method, (Wang et al., 2023) analyzes the convergence performances with linear function approximation. (Mnih et al., 2016) builds a practical system to implement the neural network approximation. In (Xie and Song, 2023), it adds KL divergence and experimentally validates the actor-critic based method with neural network approximation. For policy-based methods, in (Lan et al., 2023b), it shows simplicity compared to the other RL methods, and a linear speedup has been demonstrated in the synchronous setting. However, with an online behavior, it is not practical to perform synchronous global updates with heterogeneous computing power, and the global time consumption heavily depends on the stragglers (Mishchenko et al., 2022). This motivates us to consider the asynchronous policy-based method for FedRL. We show that our method can further reduce the time complexity compared to the synchronous FedRL approach.
2. Problem Setup
Markov decision process: We consider the Markov decision process (MDP) as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is a finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is a Markov kernel that determines transition probabilities, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a reward function, and $\gamma\in(0,1)$ is a discount factor. At each time step $t$, the agent executes an action $a_{t}\in\mathcal{A}$ from the current state $s_{t}\in\mathcal{S}$, following a stochastic policy $\pi$, i.e., $a_{t}\sim\pi(\cdot|s_{t})$. The corresponding reward is defined as $r_{t}$. The state value function is defined as
(1) $V_{\pi}(s)=\mathop{\mathbb{E}}_{a_{t}\sim\pi(\cdot|s_{t}),\atop s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right].$
Similarly, the state-action value function (Q-function) is defined as
(2) $Q_{\pi}(s,a)=\mathop{\mathbb{E}}_{a_{t}\sim\pi(\cdot|s_{t}),\atop s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,~{}a_{0}=a\right].$
An advantage function is then define as $A_{\pi}(s,a)=Q_{\pi}(s,a)-V_{\pi}(s)$. With continuous states, the policy is parameterized by $\theta\in\mathbb{R}^{d}$, and then the policy is referred as $\pi_{\theta}$ (Deep RL parameterizes $\pi_{\theta}$ by deep neural networks). A state-action visitation measure induced by $\pi_{\theta}$ is given as
(3) $\nu_{\pi_{\theta}}(s,a)=(1-\gamma)\mathop{\mathbb{E}}_{s_{0}\sim\rho}\left[\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s,~{}a_{t}=a|s_{0},~{}\pi_{\theta})\right],$
where the starting state $s_{0}$ is drawn from a distribution $\rho$. The goal of the agent is to maximize the expected discounted return defined as follows:
(4) $\max_{\theta}J(\theta)\coloneqq\mathop{\mathbb{E}}_{s\sim\rho}\left[V_{\pi_{\theta}}(s)\right].$
The gradient of $J(\theta)$ can be written as (Schulman et al., 2018):
(5) $\nabla_{\theta}J(\theta)=\mathop{\mathbb{E}}_{\tau}\left[\sum_{t=0}^{\infty}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\cdot A_{\pi_{\theta}}(s_{t},a_{t})\right],$
where $\tau=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1}\cdots)$ is a trajectory induced by policy $\pi_{\theta}$. We omit the $\theta$ notation in the gradient operation and denote the policy gradient by $g$ for short. Then, $g$ is estimated by
(6) $g(\theta,\tau)=\sum_{t=0}^{\infty}\nabla\log\pi_{\theta}(a_{t}|s_{t})\sum_{h=t}^{\infty}\gamma^{h}r(s_{h},a_{h}).$
Federated policy gradient: We aim to solve the above problem in an FL setting, where $N$ agents collaboratively train a common policy $\pi_{\theta}$. Specifically, each agent collects trajectories and corresponding reward $r(s_{h},a_{h})$ based on its local policy. Then, each agent $i$ estimates $g(\theta_{i},\tau)$ for training the model $\theta$, and the updated models are aggregated at the server. Motivated by the limitations of synchronous model aggregation in terms of time complexity, in the next section, we present our AFedPG methodology that takes an asynchronous approach to solve (4) in a FL setting.
3. Proposed Asynchronous FedPG
Figure 1. An illustration of the asynchronous federated policy gradient updates. Each agent has a local copy of the environment, and agents may collect data according to different local policies. At each iteration, the agent in the yellow color finishes the local process and then communicates with the server, while the other agents keep sampling and computing locally. In the $k$-th global iteration, $\delta_{k}\in\mathbb{N}$ is the delay, $\widetilde{\tau}_{k-\delta_{k}}$ is the sample collected according to the policy $\pi_{\widetilde{\theta}_{k-\delta_{k}}}$, and $d_{k-\delta_{k}}$ is the updating direction calculated from the sample $\widetilde{\tau}_{k-\delta_{k}}$.
The proposed algorithm consists of $K$ global iterations, indexed by $k=0,1,\dots,K-1$. We first introduce the definition of concurrency and delay in our asynchronous federated setting. We then present the proposed AFedPG methodology.
Definition 3.1 ().
(Concurrency) We denote $\mathcal{C}_{k}$ as the set of active agents in the $k$-th global iteration, and define $\omega_{k}\coloneqq|\mathcal{C}_{k}|$ as the concurrency. We define the average and maximum concurrency as $\bar{\omega}\coloneqq\frac{1}{K}\sum_{k=0}^{K-1}\omega_{k}$, and $\omega_{\max}\coloneqq\max\omega_{k}$, respectively.
In each global iteration of AFedPG, the server applies only one gradient to update the model from the agent who has finished its local computation, while the other $N-1$ agents keep computing local gradients (unapplied gradients) in parallel. In this asynchronous setting, the models used in each agent are outdated, as the server keeps updating the model. Thus, introduce the notion of delay (or staleness) $\delta\in\mathbb{N}$, which measures the difference between the current global iteration and the past global iteration when the agent received the updated model from the server.
Definition 3.2 ().
(Delay) In the $k$-th global iteration, we denote the delay of the applied gradient as $\delta_{k}\in\mathbb{N}$, and the delays of the unapplied gradients as $\{\delta_{k}^{i}\}_{i\in\mathcal{C}_{k}\setminus\{j_{k}\}}$, where $j_{k}$ denotes the agent that communicates with the server. $\delta_{k}^{i}$ is the difference between the $k$-th iteration and the iteration where the agent $i$ started to compute the latest gradient. In the final $K$-th global iteration, we have $K$ applied gradients $\{\delta_{k}\}_{k=0}^{K-1}$ and unapplied gradients $\{\delta_{k}^{i}\}_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}$. The average and maximum delays can be expressed as
(7) $\bar{\delta}~{}\coloneqq~{}\frac{1}{K-1+|\mathcal{C}_{K}|}\big{(}\sum_{k=0}^{K-1}\delta_{k}+\sum_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}\delta_{K}^{i}\big{)},$
and $\delta_{\max}\coloneqq\max(\max_{k}\delta_{k},\max_{i\in\mathcal{C}_{K}}\delta_{K}^{i})$, respectively.
Asynchronous FedPG: Federated learning is a paradigm in which $N$ agents collaboratively train a common global policy with parameters $\theta\in\mathbb{R}^{d}$. As shown in Figure 1, during the training process, agent $i$ collects trajectories in an online behavior, and computes gradients or updating directions using its local trajectories (also known as samples). Then, agent $i$ transmits local gradients to the central server. In particular, in the $k$-th global iteration, AFedPG training consists of the following three steps:
  • Local computation and uplink transmission: Agent $i$ receives the previous policy $\pi_{\theta_{j}}$ from the server in the $j$-th global iteration. Agent $i$ collects its own local trajectory $\tau_{j}$ based on its current policy $\pi_{\theta_{j}}$. Agent $i$ then computes its local updating direction $d_{j}\in\mathbb{R}^{d}$ based on the trajectory $\tau_{j}$, and sends it back to the server.
  • Server-side model update: The server starts to operate the $k$-th global iteration as soon as it receives $d_{j}$ from agent $i$. Thus, denote $d_{k-\delta_{k}}=d_{j}$, where $\delta_{k}$ is the delay in the $k$-th global iteration. The server updates global policy parameters by $\theta\leftarrow\theta-\eta d_{k-\delta_{k}}$, where $\eta$ is the learning rate.
  • Downlink transmission: The server transmits the current global policy parameters $\theta\in\mathbb{R}^{d}$ back to agent $i$ as soon as it finishes the global update.
The server side procedure of AFedPG is shown in Algorithm 1 and the process at the agent side is shown in Algorithm 2. In the $k$-th iteration, the server operates one global update (Steps 4 and 5) as soon as it receives a direction $d_{k-\delta_{k}}$ from an agent with delay $\delta_{k}$. After the global update, the server sends the updated model back to that agent. In Algorithm 2, after receiving the global model from the server, an agent first gets model parameters $\widetilde{\theta}_{k}$ according to Step 2, and then collects samples based on the policy $\pi_{\widetilde{\theta}_{k}}$ (Steps 3). At last, the agent computes the updating direction $d_{k}$, and sends it to the server as soon as it finishes the local process. Overall, all agents conduct local computation in parallel, but the global model is updated in an asynchronous manner as summarized in Figure 1.
Delay-adaptive lookahead and normalized update: In Step 5 of Algorithm 1, to handle the updates with various delays, we use normalized gradients with controllable sizes. Specifically, the error term $\|e_{k}\|$ in Lemma 4.6 is related to $\|\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\|$ and $\|\nabla J(\theta_{k-1})-\nabla J(\theta_{k})\|$. With the smoothness in Lemma 4.4, we are able to bound the error as $\|\theta_{k-1}-\theta_{k}\|=\eta_{k-1}$. The details of the boundaries are shown in (34).
In Step 2 of Algorithm 2 (in blue), we design a delay-adaptive lookahead update, i.e.,
(8) ${\theta}_{k}=(1-\alpha_{k-\delta_{k}}){\theta}_{k-1}+\alpha_{k-\delta_{k}}\widetilde{\theta}_{k}.$
Notably, this is not the momentum method, because the orders are different. The agent collects samples according to $\widetilde{\theta}_{k}$, which only happens in RL. Thus, it is fundamentally different from conventional FL. This mechanism cancels out the second-order Hessian correction terms:
(9) $(1-\alpha_{k-\delta_{k}})\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})+\alpha_{k-\delta_{k}}\nabla^{2}J(\theta_{k})(\widetilde{\theta}_{k}-\theta_{k})\rightarrow 0,$
and thus assists the convergence analysis. The details are shown in Appendix A.2.
Algorithm 1 AFedPG: Server.
0:  MDP $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$; Number of iterations $K$; Step size $\eta_{k}$, $\alpha_{k}$; Initial $\theta_{0}\in\mathbb{R}^{d}$.
1:  Broadcast $\theta_{0}$ to $N$ agents.
2:  for $k=0,\cdots,K$ do
[…]
4:          Receive $d_{k-\delta_{k}}$ from the agent $i_{k}$.
5:          $\theta_{k+1}\leftarrow\theta_{k}+\eta_{k}\frac{d_{k-\delta_{k}}}{\|d_{k-\delta_{k}}\|}$
[…]
7:          Transmit $\theta_{k+1}$ back to the agent $i_{k}$.
8:  end for
8:  ${\theta^{K+1}}$
Algorithm 2 AFedPG: Agent $i$ Update, $i=1,\cdots,N$.
0:  $\theta_{k}\in\mathbb{R}^{d}$
1:  Receive $\theta_{k}$ from the server.
2:  $\widetilde{\theta}_{k}\leftarrow\theta_{k}+\frac{1-\alpha_{k-\delta_{k}}}{\alpha_{k-\delta_{k}}}(\theta_{k}-\theta_{k-1})$
3:  $\widetilde{\tau}_{k}\sim p(\cdot|\pi_{\widetilde{\theta}_{k}})$
4:  $d_{k}\leftarrow(1-\alpha_{k})d_{k-1}+\alpha_{k}g(\widetilde{\tau}_{k},\widetilde{\theta}_{k})$
5:  When finish computing, transmit ${d}_{k}$ to the server.
5:  ${d}_{k}$
4. Convergence Analysis
In this section, we derive the convergence rate of AFedPG with the following criteria for the global convergence.
Criteria: In this work, we focus on the global convergence, i.e., finding the parameter $\theta$ s.t. $J^{\star}-J(\theta)\leq\epsilon^{\prime}$, where $J^{\star}$ is the optimal expected return.
4.1. Assumptions
In order to derive the global convergence rates, we make the following standard assumptions (Agarwal et al., 2021b; Ding et al., 2020; Liu et al., 2020; Papini et al., 2018; Xu et al., 2020) on policy gradients, and rewards.
Assumption 4.1 ().
  1. (1)
[…]
  2. (2)
    The score function is $M_{h}$-Lipschitz continuous. In other words, $\forall~{}\theta_{i},~{}\theta_{j}\in\mathbb{R}^{d},~{}s\in\mathcal{S},~{}\text{and}~{}a\in\mathcal{A}$, we have
    (10) $\left\|\nabla\log\pi_{\theta_{i}}(a\mid s)-\nabla\log\pi_{\theta_{j}}(a\mid s)\right\|$ $\leq M_{h}\left\|\theta_{i}-\theta_{j}\right\|.$
  3. (3)
    The reward function is bounded as $r(s,a)\in[0,R],\text{ for all }s\in\mathcal{S}$, and $a\in\mathcal{A}$.
These standard assumptions naturally state the reward, first-order and second-order derivatives of the score function are not infinite, and they hold with common practical parametrization methods, e.g., softmax policies. With function approximation $\pi_{\theta}$, the approximation error may not be 00 in practice. Thus, we follow previous works (Liu et al., 2020; Agarwal et al., 2021b; Ding et al., 2022; Huang et al., 2020) with an assumption on the expressivity of the policy parameterization class.
Assumption 4.2 ().
(Function approximation) $\exists~{}\epsilon_{bias}\geq 0$ s.t. for $\forall~{}\theta\in\mathbb{R}^{d}$, the transfer error satisfies
(11) $\mathbb{E}[\big{(}A_{\pi_{\theta}}(s,a)-(1-\gamma)u^{\star}(\theta)^{\top}\nabla\log\pi_{\theta}(a\mid s)\big{)}^{2}]\leq\epsilon_{bias},$
where $u^{\star}(\theta)\coloneqq F_{\rho}(\theta)^{\dagger}\nabla J(\theta)$, and $F_{\rho}(\theta)^{\dagger}$ is the Moore-Penrose pseudo-inverse of the Fisher matrix $F_{\rho}$.
It means that the parameterized policy $\pi_{\theta}$ makes the advantage function $A_{\pi_{\theta}}(s,a)$ approximated by the score function $\nabla\log\pi_{\theta}(a\mid s)$ as the features. This assumption is widely used with Fisher-non-degenerate parameterization. $\epsilon_{bias}$ can be very small with rich neural network parameterization (Wang et al., 2020), and 00 with a soft-max parameterization (Ding et al., 2022).
4.2. Convergence Rates
Based on the above assumptions, the converges rate of AFedPG is given in Theorem 4.3.
Theorem 4.3.
Let Assumption 4.1 and 4.2 hold. With suitable learning rates $\eta_{k}$ and $\alpha_{k}$, after $K$ global iterations, AFedPG satisfy
(12) $J^{\star}-\mathbb{E}[J(\theta_{K})]~{}\leq~{}\mathcal{O}\big{(}{K^{-\frac{2}{5}}}\cdot(1-\gamma)^{-3}\big{)}+\frac{\sqrt{\epsilon_{bias}}}{1-\gamma},$
where $\epsilon_{bias}$ is from (11). Thus, to satisfy $J^{\star}-J(\theta_{K})\leq\epsilon+\frac{\sqrt{\epsilon_{bias}}}{1-\gamma}$, we need $K=\mathcal{O}(\frac{\epsilon^{-2.5}}{(1-\gamma)^{7.5}})$ iterations. As only one trajectory is required in each iteration, the number of trajectories is equal to $K$, i.e., the sample complexity is $\mathcal{O}(\frac{\epsilon^{-2.5}}{(1-\gamma)^{7.5}})$.
Comparison to the synchronous setting: Synchronous FedPG needs $\mathcal{O}(\epsilon^{-2.5})$ trajectories in total, and each agent needs $\mathcal{O}(\frac{\epsilon^{-2.5}}{N})$ trajectories. However, in synchronous FedGP the server has to wait for the slowest agent at each global step, which can still slow down the training process. Let $t_{i}$ denote the time consumption for agent $i=1,\cdots,N$ at local steps with finite values. Then, the waiting time on the server for each step becomes $t_{\max}\coloneqq\max t_{i}$ for FedPG. Our AFedPG approach keeps the same sample complexity as FedPG, but the server processes the global step as soon as it receives an update, speeding up training. Specifically, the average waiting time on the server is $\bar{t}\coloneqq\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}}<\frac{t_{\max}}{N}$ at each step. Thus, the asynchronous FedPG achieves better time complexity than the synchronous approach regardless of the delay pattern. The advantage is significant when $t_{\max}\gg t_{min}$, which occurs in many practical settings with heterogeneous computation powers across different agents. We illustrate the advantage of AFedPG over synchronous FedPG in Figure 2. As the server only operates one simple summation, without loss of generality, the time consumption at the server-side is negligible.
Figure 2. Comparison of time consumptions between synchronous and asynchronous approaches. The circled numbers denote the indices of global steps.
4.3. Key Lemmas
We list some useful lemmas to construct the proof sketch of Theorem 4.3. The details of the proof can be found in Appendix A.
Under Assumption 4.1 on score functions, the following lemma holds based on the results in (Zhang et al., 2020).
Lemma 4.4.
The gradient of the expected return is $L_{g}$-continuous and $L_{h}$-smooth as follows
(13) $\|\nabla J(\theta)-\nabla J(\theta^{\prime})\|~{}\leq$ $~{}L_{g}\|\theta-\theta^{\prime}\|,$
$\|\nabla^{2}J(\theta)-\nabla^{2}J(\theta^{\prime})\|~{}\leq$ $~{}L_{h}\|\theta-\theta^{\prime}\|,$
where $L_{g}\coloneqq\frac{R(M_{g}^{2}+M_{h})}{(1-\gamma)^{2}}$ and $L_{h}\coloneqq\frac{RM_{g}^{3}(1+\gamma)}{(1-\gamma)^{3}}+\frac{RM_{g}M_{h}}{(1-\gamma)^{2}}+\mathcal{O}\big{(}(1-\gamma)^{-1}\big{)}$.
Under Assumption 4.1 and 4.2, we utilize the result in (Ding et al., 2022) as the Lemma 4.5.
Lemma 4.5.
(Relaxed weak gradient domination) Under Assumptions 4.1 and 4.2, it holds that
(14) $\|\nabla J(\theta)\|+\epsilon_{g}~{}\geq~{}\sqrt{2\mu}(J^{\star}-J(\theta)),$
where $\epsilon_{g}=\frac{\mu_{F}\sqrt{\epsilon_{bias}}}{M_{g}(1-\gamma)}$ and $\mu=\frac{\mu_{F}^{2}}{2M_{g}^{2}}$.
Based on Lemma 4.4, we derive our milestone, the ascent lemma with delayed updates, as follows:
Lemma 4.6.
(Ascent Lemma with Delay) Under Assumptions 4.1, it holds that
(15) $-J(\theta_{k+1})~{}\leq~{}-J(\theta_{k})-\frac{1}{3}\eta_{k}\|\nabla J(\theta_{k})\|+\frac{8}{3}\eta_{k}\|e_{k}\|+\frac{L_{g}}{2}\eta_{k}^{2},$
where $e_{k}\coloneqq d_{k-\delta_{k}}-\nabla J(\theta_{k})$.
This ascent lemma is specific to the asynchronous setting, which constructs the lower bound for the global increment, $J(\theta_{k+1})-J(\theta_{k})$, through the normalized policy gradients and our updating rules with delay $\delta_{k}$.
Table 2. Description of the MuJoCo environments.
Tasks Agents Action Space Dimension State Space Dimension
Swimmer-v4 Three-link swimming robot $2$ $8$
Hopper-v4 Two-dimensional one-legged robot $3$ $11$
Humanoid-v4 Three-dimensional bipedal robot $17$ $376$
(a) Swimmer-v4
(b) Hopper-v4
(c) Humanoid-v4
Figure 3. Virtualization of the three tasks in the MuJoCo environments.
Proof.
With the smoothness of the expected return $J(\theta)$ and the updating rule, we have
(16) $-J(\theta_{k+1})~{}$ $\leq~{}-J(\theta_{k})-\langle\nabla J(\theta_{k}),\theta_{k+1}-\theta_{k}\rangle+\frac{L_{g}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$=~{}-J(\theta_{k})-\eta_{k}\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}+\frac{L_{g}}{2}\eta_{k}^{2}.$
If $\|e_{k}\|\leq\frac{1}{2}\|\nabla J(\theta_{k})\|$, we have
(17) $-\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}~{}$ $=~{}-\frac{\|\nabla J(\theta_{k})\|^{2}+\langle\nabla J(\theta_{k}),e_{k}\rangle}{\|d_{k-\delta_{k}}\|}$
$\leq~{}\frac{-\|\nabla J(\theta_{k})\|^{2}+\|\nabla J(\theta_{k})\|\|e_{k}\|}{\|d_{k-\delta_{k}}\|}$
$\leq~{}\frac{-\|\nabla J(\theta_{k})\|^{2}+\frac{1}{2}\|\nabla J(\theta_{k})\|^{2}}{\|\nabla J(\theta_{k})\|+\|e_{k}\|}\leq~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|.$
If $\|e_{k}\|\geq\frac{1}{2}\|\nabla J(\theta_{k})\|$, we have
(18) $-\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}\leq~{}\|\nabla J(\theta_{k})\|$
$=~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|+\frac{4}{3}\|\nabla J(\theta_{k})\|\leq~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|+\frac{8}{3}\|e_{k}\|.$
Combine these two conditions, and plugging the result in (16), the lemma can be proved as
(19) $-J(\theta_{k+1})~{}$ $\leq~{}-J(\theta_{k})-\eta_{k}\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}+\frac{L_{g}}{2}\eta_{k}^{2}$
$\leq~{}-J(\theta_{k})-\frac{1}{3}\eta_{k}\|\nabla J(\theta_{k})\|+\frac{8}{3}\eta_{k}\|e_{k}\|+\frac{L_{g}}{2}\eta_{k}^{2}.$
Next, we construct the relationship between the average concurrency $\bar{\omega}$ and the average delay $\bar{\delta}$ in Lemma 4.7. This lemma gives the boundary (ours) of delays as a corollary of the result in (Koloskova et al., 2022).
Lemma 4.7.
The average delay $\bar{\delta}$ depends on the average concurrency $\bar{\omega}$ as
(20) $\bar{\delta}=\frac{K+1}{K-1+|\mathcal{C}_{K}|}\bar{\omega}\leq\bar{\omega}\leq N.$
Proof.
Recall that $\{\delta_{k}^{i}\}_{i\in\mathcal{C}_{k}\setminus\{j_{k}\}}$ is the set of delays at the $k$-th global steps. After one global step, the number of cumulative delays over all agents increases by the current concurrency. Thus, we have the following connection
(21) $\sum_{i=0}^{k}\delta_{i}+\sum_{i\in\mathcal{C}_{k+1}\setminus\{j_{k+1}\}}\delta^{i}_{k+1}~{}$ $=~{}\sum_{i=0}^{k-1}\delta_{i}+\sum_{i\in\mathcal{C}_{k}\setminus\{j_{k}\}}\delta^{i}_{k}+\omega_{k+1}.$
At the zeroth step, there is no delay. Therefore, we have $\delta^{i}_{0}=0$ for all agents. Unrolling the above expression, at the $K$-th step, we have
(22) $\sum_{i=0}^{K-1}\delta_{i}+\sum_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}\delta^{i}_{K}~{}$ $=~{}\sum_{i=0}^{K}\omega_{k+1}~{}=~{}(K+1)\bar{\omega}.$
According to (7) and $|\mathcal{C}_{K}|\geq 2$, we achieve
(23) $\bar{\delta}~{}=~{}\frac{K+1}{K-1+|\mathcal{C}_{K}|}\bar{\omega}\leq\bar{\omega}\leq N.$
In practice, we need to use all resources, and thus, $\bar{\omega}=\omega_{\max}=N$. Since the maximum concurrency is equal to the number of agents $N$, we have the boundary of the average delay as $\bar{\delta}\leq\omega_{\max}\leq N$. ∎
4.4. Proof sketch of Theorem 4.3
Here, we briefly give the main idea of our proof of Theorem 4.3. The complete details and the technical lemmas are provided in Appendix A.
First, we prove the ascent lemma 4.6 based on Lemma 4.4 and the updating rule of AFedPG in Algorithm 1 and 2. We then plug it into (14) in Lemma 4.5 in the expectation form, and get the following inequality
(24) $J^{\star}-\mathbb{E}[J(\theta_{k+1})]~{}\leq$ $~{}(1-\frac{\sqrt{2\mu}\eta_{k}}{3})\big{(}J^{\star}-\mathbb{E}[J(\theta_{k})]\big{)}$
$~{}+\frac{\eta_{k}}{3}\epsilon_{g}+\frac{8}{3}\eta_{k}\mathbb{E}[\|e_{k}\|]+\frac{L_{g}}{2}\eta_{k}^{2}.$
Second, we bound the expected error term $\mathbb{E}[\|e_{K}\|]$ in the ascent lemma with delayed updates. Here we utilize the normalization and look-ahead sampling in Algorithm 1 and 2 to bound the error with an average delay term $\bar{\delta}$ from Lemma 4.7. With suitable learning rates as $\alpha_{k}=(\frac{1}{k+1})^{\frac{4}{5}}$ and $\eta_{k}=\eta_{0}\frac{1}{k+1}$, the upper bound of $\mathbb{E}[\|e_{K}\|]$ is given as follows
(25) $\mathbb{E}[\|e_{K}\|]{\leq}$ $~{}\alpha_{K}\sigma_{g}+c_{1}\sqrt{\alpha_{K}}\sigma_{g}+\frac{9}{4}c_{2}L_{h}\frac{\eta_{K}^{2}}{\alpha_{K}^{2}}+2c_{3}L_{g}\frac{\eta_{K}^{2}}{\alpha_{K}}\bar{\delta},$
where $c_{1}$, $c_{2}$ and $c_{3}$ are constants, and their values are given in Appendix A.2.
Third, unroll the recursion of (24), and then plug the upper bound result (25) into the unrolled form. We now have the following inequality
(26) $J^{\star}$ $-\mathbb{E}[J(\theta_{K})]~{}\leq~{}\frac{\eta_{0}}{3}\epsilon_{g}+\frac{J^{\star}-\mathbb{E}[J(\theta_{0})]}{(K+1)^{2}}+c_{3}\frac{16\eta_{0}^{2}}{3}\frac{L_{g}}{(K+1)^{\frac{7}{5}}}\bar{\delta}$
$+\frac{\eta_{0}^{2}}{2}\frac{L_{g}}{K+1}+\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{4}{5}}}+c_{1}\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{2}{5}}}+6c_{2}\eta_{0}^{3}\frac{L_{h}}{(K+1)^{\frac{2}{5}}}.$
With an initial learning rate as $\eta_{0}=\frac{3\mu_{F}}{M_{g}}$, we achieve the global convergence rate in Theorem 4.3.
5. Experiments
5.1. Setup
Environment: To validate the effectiveness of our approach via experiments, we consider three popular MuJoCo environments (Swimmer-v4, Hopper-v4, and Humanoid-v4) (Todorov et al., 2012) with the MIT License. Environmental details are described in Table 2 and the tasks are visualized in Figure 3.
Table 3. Hyperparameter and MLP settings.
Hyperparameter Setting
Task Swimmer-v4 Hopper-v4 Walker2D-v4 Humanoid-v4
MLP $64\times 64$ $256\times 256$ $256\times 256\times 256$ $512\times 512\times 512$
Activation function ReLU ReLU ReLU ReLU
Output function Tanh Tanh Tanh Tanh
Learning rate ($\alpha$) $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Discount ($\gamma$) $0.99$ $0.99$ $0.99$ $0.99$
Timesteps ($T$) $2048$ $1024$ $512$ $512$
Iterations ($K$) $1\times 10^{3}$ $2\times 10^{3}$ $3\times 10^{3}$ $1.5\times 10^{4}$
Learning rate ($\eta$) $3\times 10^{-4}$ $1\times 10^{-4}$ $5\times 10^{-5}$ $1\times 10^{-5}$
(a) Swimmer-v4
(b) Hopper-v4
(c) Humanoid-v4
Figure 4. Reward performances of AFedPG ($N=2,4,8$) and PG ($N=1$) on various MoJuCo environments, where $N$ is the number of federated agents. The shadowed areas are confidence intervals with $95\%$ confidence level.
Measurement: All convergence performances are measured over $10$ runs with random seeds from 00 to $9$. The solid lines in Figure 4 and 5 are the averaged results, and the shadowed areas are confidence intervals with the confidence level $95\%$. The lines are smoothed for better visualization.
Implementation: Policies are parameterized by fully connected multi-layer perceptions (MLPs) with settings listed in Table 3. We follow the practical settings in stable-baselines3 (Raffin et al., 2021) to update models with generalized advantage estimation (GAE) ($0.95$) (Schulman et al., 2018) in our implementation. We use PyTorch (Paszke et al., 2019) to implement deep neural networks. All tasks are trained on NVIDIA A100 GPUs with $40$ GB of memory.
Performance metrics: We consider performance metrics as follows:
  1. (1)
    Rewards: the average trajectory rewards collected at each iteration;
  2. (2)
    Convergence: rewards versus iterations during the training process;
  3. (3)
    Time consumption: global time with certain numbers of collected samples.
5.2. Results
Sample complexity: First, to verify the improvement of sample complexity in the first row of Table 1, we evaluate the speedup effects of the number of federated agents $N$. In Figure 4, with different numbers of agents, we test the convergence performances of AFedPG ($N=2,4,8$) and the single agent PG ($N=1$). The x-axis is the number of samples collected by each agent on average, and the y-axis is the reward. In all three tasks, AFedPG beats PG: AFedPG converges faster, has lower variances, and achieves higher final rewards when more agents are involved in collecting trajectories and estimating policy gradients.
(a) Swimmer-v4
(b) Hopper-v4
(c) Humanoid-v4
Figure 5. Global time of AFedPG and FedPG with certain numbers of collected samples on various MoJuCo environments, where $N$ is the number of federated agents. The shadowed areas are confidence intervals with $95\%$ confidence level.
Figure 6. The time complexity speedup of AFedPG compared to FedPG. The x-axis is the heterogeneous ratio of computing power measured by $\frac{t_{\max}}{t_{\min}}$.
Global time complexity: Second, to verify the improvement of global time complexity in the second row of Table 1, we compared the time consumption in the asynchronous and synchronous settings. In Figure 5, we set $N=4,8$ and fix the number of samples collected by all agents. Here, $t_{\max}$ is about $4$ times more than $t_{\min}$. The numbers of total trajectories are $8\times 10^{3}$, $16\times 10^{3}$, and $8\times 10^{4}$ for the Swimmer task, Hopper task, and Humanoid task individually when $N=8$. When $N=4$, the numbers are halved. In all three environments, AFedPG has much lower time consumption compared to the synchronous FedPG.
Effect of computation heterogeneity: At last, we study the effect of the computation heterogeneity among federated agents. The heterogeneity of computing powers is measured by the ratio $\frac{t_{\max}}{t_{\min}}$, and the effect is measured by the speedup, which is the global time of FedPG divided by that of AFedPG. Without loss of generality, we test the performances with one dragger here, which is one agent has $t_{\max}$ time consumption, and all the other agents have $t_{\min}$ time consumption. In Figure 6, the results show that the speedup increases in a nearly linear behavior as the heterogeneity ratio $\frac{t_{\max}}{t_{\min}}$ increases. On the other hand, the more agents that participate in the training process, the higher speedup they achieve. Therefore, AFedPG has the potential to scale up into a very large mobile system, especially when there are some extreme stragglers.
6. Conclusion
In this paper, we proposed AFedPG, a novel asynchronous federated reinforcement learning framework, which updates the global model using policy gradients from $N$ agents. To handle the challenge of lagged policies in the asynchronous setting, we designed a delay-adaptive lookahead technique and used normalized updates to integrate policy gradients. With the above methods, AFedPG achieves speedup for both sample complexity and time complexity. First, our AFedPG method achieves $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ sample complexity at each agent. Compared to the single agent setting (PG) with $\mathcal{O}(\epsilon^{-2.5})$ sample complexity, it enjoys a linear speedup with respect to the number of agents $N$. Second, compared to synchronous FedPG, AFedPG improves the time consumption from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$, where $t_{i}$ denotes the time consumption in each iteration at the agent $i$, and $t_{\max}$ is the largest one. The latter complexity $\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$ is always smaller than the former one, and this improvement can be significant in a large-scale federated setting with heterogeneous computing powers ($t_{\max}\gg t_{\min}$). Finally, we empirically verified the performances of AFedPG compared to PG and FedPG in three MuJoCo environments with different $N$. We also demonstrate the improvements with different computing heterogeneity. In summary, AFedPG achieves speedup in terms of both sample complexity and time complexity when it is scaled up to a large mobile system, especially with computing power heterogeneity.
Appendix A Proofs
A.1. Technical Lemmas
Lemma A.1.
For arbitrary $n$ vectors $\{\mathbf{a}_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$, we have
(27) $\|\sum_{i=1}^{n}\mathbf{a}_{i}\|^{2}~{}\leq~{}n\sum_{i=1}^{n}\|\mathbf{a}_{i}\|^{2}.$
Lemma A.2.
Let $\alpha_{k}=(\frac{c}{t+c})^{p}$. $\forall p\in[0,1]$ and $c\geq 1$, we have
(28) $1-\alpha_{k+1}~{}\leq~{}\frac{\alpha_{k+1}}{\alpha_{k}}.$
Proof.
(29) $1-\alpha_{k+1}~{}=$ $~{}1-(\frac{c}{t+1+c})^{p}$
[…]
Straightforwardly, we have
(30) $\prod_{i=k}^{K-1}(1-\alpha_{i+1})~{}\leq$ $~{}\frac{\alpha_{K}}{\alpha_{k}}.$
Lemma A.3.
Let $\alpha_{k}=(\frac{1}{k+1})^{p}$ and $\eta_{k}=\eta_{0}(\frac{1}{k+1})^{q}$. $\forall p\in[0,1)$, $q\geq 0$ and $\eta_{0}\geq 0$, we have
(31) $\sum_{k=0}^{K-1}\eta_{k}\prod_{i=k+1}^{K-1}(1-\alpha_{i})~{}\leq~{}c(p,q)\frac{\eta_{K}}{\alpha_{K}},$
where $c(p,q)\coloneqq\frac{2^{q-p}}{1-p}a\exp{\big{(}(1-p)2^{p}a^{1-p}\big{)}}$ is a constant, and $a=\max\big{(}(\frac{q}{(1-p)2^{p}})^{\frac{1}{1-p}},(\frac{2(q-p)}{(1-p)^{2}})^{\frac{1}{1-p}}\big{)}$.
A.2. Proof of Theorem 4.3
Denote $\xi_{k}\coloneqq g(\widetilde{\tau}_{k},\widetilde{\theta}_{k})-\nabla J(\widetilde{\theta}_{k})$. First, according to the updating rules in Algo. 1 and Algo. 2 , we expand the error term in Lemma 4.6 as follows
(32) $e_{k}~{}=$ $~{}(1-\alpha_{k-\delta_{k}})d_{\delta_{k-1}}+(1-\alpha_{k-\delta_{k}})\big{(}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})\big{)}+\alpha_{k-\delta_{k}}\big{(}g(\widetilde{\tau}_{k-\delta_{k}},\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\theta_{k})\big{)}$
[…]
$+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\big{)}+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k})-\nabla J(\theta_{k})\big{)}$
[…]
$+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\big{)}+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}-(1-\alpha_{k-\delta_{k}})\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})-\alpha_{k-\delta_{k}}\nabla^{2}J(\theta_{k})(\widetilde{\theta}_{k}-\theta_{k})}$
${{=}}$ $~{}\alpha_{k-\delta_{k}}\xi_{k-\delta_{k}}+(1-\alpha_{k-\delta_{k}})e_{k-1}+(1-\alpha_{k-\delta_{k}})\big{(}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}$
$+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\big{)}+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}.$
Remark: The Hessian correction terms (in blue) are equal to 00 according to our updating rules (delay-adaptive lookahead update) in Step 2 of Algo. 2. We design this technique to cancel out the second-order terms, and thus achieve the desired convergence rate.
Next, we denote
(33) $A_{k}~{}\coloneqq$ $~{}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k}),$
$B_{k}~{}\coloneqq$ $~{}\nabla J(\widetilde{\theta}_{k})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k}),$
$C_{k}~{}\coloneqq$ $~{}\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k}).$
With the smoothness in Lemma 4.4 and the non-increasing learning rates, we have
(34) $\|A_{k}\|~{}$ $\leq~{}L_{h}\|\theta_{k-1}-\theta_{k}\|^{2}=L_{h}\eta_{k-1}^{2},$
$\|B_{k}\|~{}$ $\leq~{}L_{h}\|\widetilde{\theta}_{k}-\theta_{k}\|^{2}=L_{h}\frac{(1-\alpha_{k-\delta_{k}})^{2}}{\alpha_{k-\delta_{k}}^{2}}\eta_{k-1}^{2},$
$\|C_{k}\|~{}$ $\leq~{}L_{g}\|\widetilde{\theta}_{k-\delta_{k}}-\widetilde{\theta}_{k}\|$
$=~{}L_{g}\left\|\sum_{i=k-\delta_{k}}^{k-1}\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|$
$\leq~{}L_{g}\sum_{i=k-\delta_{k}}^{k-1}\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\|$
$\leq~{}L_{g}\sum_{i=k-\delta_{k}}^{k-1}(\frac{1}{\alpha_{i+1-\delta_{i+1}}}\eta_{i}+\frac{1-\alpha_{i-\delta_{i}}}{\alpha_{i-\delta_{i}}}\eta_{i-1})$
$\leq~{}L_{g}\sum_{i=k-\delta_{k}}^{k-1}(\frac{2}{\alpha_{k-1}}\eta_{k-\delta_{k}})$
$=~{}\delta_{k}L_{g}\frac{2\eta_{k-\delta_{k}}}{\alpha_{k-1}}.$
Denote $\beta_{k}\coloneqq\prod_{i=k+1}^{K}(1-\alpha_{i-\delta_{i}})$ with $\beta_{K}=1$. Unrolling the recursion (32), we have
(35) $e_{K}~{}$ $=~{}\beta_{0}e_{0}+\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\xi_{k-\delta_{k}}+\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}A_{k}+\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}B_{k}.$
Choose learning rates $\alpha_{k}=(\frac{1}{k+1})^{\frac{4}{5}}$ and $\eta_{k}=\eta_{0}\frac{1}{k+1}$, where $\eta_{0}$ is a constant and we will show the value later. Using Jensen’s inequality and technical lemmas, we achieve the following bound
(36) $\mathbb{E}[\|e_{K}\|]~{}\leq$ $~{}\beta_{0}\mathbb{E}[\|e_{0}\|]+\mathbb{E}\left[\left\|\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\xi_{k-\delta_{k}}\right\|^{2}\right]^{\frac{1}{2}}+\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}\mathbb{E}[\|A_{k}\|]$
$+\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}(\mathbb{E}[\|B_{k}\|]+\mathbb{E}[\|C_{k}\|])$
(${{\leq}}$ $~{}\beta_{0}\sigma_{g}+\left(\sum_{k=1}^{K}\alpha_{k-\delta_{k}}^{2}\beta_{k}^{2}\mathbb{E}\left[\left\|\xi_{k-\delta_{k}}\right\|^{2}\right]\right)^{\frac{1}{2}}+L_{h}\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}\eta_{k-1}^{2}$
$+L_{h}\sum_{k=1}^{K}\beta_{k}\frac{(1-\alpha_{k-\delta_{k}})^{2}}{\alpha_{k-\delta_{k}}}\eta_{k-1}^{2}+2L_{g}\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\delta_{k}\frac{\eta_{k-\delta_{k}}}{\alpha_{k-1}}$
[…]
(${{\leq}}$ $~{}\alpha_{K}\sigma_{g}+\left(\sum_{k=1}^{K}\alpha_{k-\delta_{k}}^{2}\beta_{k}^{2}\right)^{\frac{1}{2}}\sigma_{g}+\frac{9L_{h}}{4}\sum_{k=1}^{K}\frac{\eta_{k}^{2}}{\alpha_{k}}\beta_{k}+2L_{g}\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\delta_{k}\frac{\eta_{k-\delta_{k}}}{\alpha_{k-1}}$
(${{\leq}}$ $~{}\alpha_{K}\sigma_{g}+c_{1}\sqrt{\alpha_{K}}\sigma_{g}+\frac{9}{4}c_{2}L_{h}\frac{\eta_{K}^{2}}{\alpha_{K}^{2}}+2c_{3}L_{g}\frac{\eta_{K}^{2}}{\alpha_{K}}\bar{\delta},$
where $c_{1}\coloneqq\sqrt{c(\frac{4}{5},\frac{8}{5})}$, $c_{2}\coloneqq c(\frac{4}{5},\frac{6}{5})$, $c_{3}\coloneqq c(\frac{4}{5},0)$ are constants, and the values of $c(\cdot,\cdot)$ are defined in Lemma A.3. After plugging the result into Lemma 4.6 and using Lemma 4.5, we achieve
(37) $J^{\star}-\mathbb{E}[J(\theta_{k+1})]~{}\leq$ $~{}(1-\frac{\sqrt{2\mu}\eta_{k}}{3})\big{(}J^{\star}-\mathbb{E}[J(\theta_{k})]\big{)}+\frac{\eta_{k}}{3}\epsilon_{g}+\frac{8}{3}\eta_{k}\mathbb{E}[\|e_{k}\|]+\frac{L_{g}}{2}\eta_{k}^{2}$
(${{\leq}}$ $~{}(1-\frac{\sqrt{2\mu}\eta_{k}}{3})\big{(}J^{\star}-\mathbb{E}[J(\theta_{k})]\big{)}+\frac{\eta_{k}}{3}\epsilon_{g}+\frac{L_{g}}{2}\eta_{k}^{2}$
$+\frac{8}{3}\eta_{k}\big{(}\alpha_{k}\sigma_{g}+c_{1}\sqrt{\alpha_{k}}\sigma_{g}+\frac{9}{4}c_{2}L_{h}\frac{\eta_{k}^{2}}{\alpha_{k}^{2}}+2c_{3}L_{g}\frac{\eta_{k}^{2}}{\alpha_{k}}\bar{\delta}\big{)}.$
Unrolling this recursion, we have
(38) $J^{\star}-\mathbb{E}[J(\theta_{K})]~{}\leq$ $~{}\frac{\eta_{0}}{3}\epsilon_{g}+\frac{J^{\star}-\mathbb{E}[J(\theta_{0})]}{(K+1)^{2}}+c_{3}\frac{16\eta_{0}^{2}}{3}\frac{L_{g}}{(K+1)^{\frac{7}{5}}}\bar{\delta}+\frac{\eta_{0}^{2}}{2}\frac{L_{g}}{K+1}+\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{4}{5}}}$
$+c_{1}\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{2}{5}}}+6c_{2}\eta_{0}^{3}\frac{L_{h}}{(K+1)^{\frac{2}{5}}}.$
Note that, introduced in Lemma 4.4, the discount factor $\gamma$ is contained in $L_{g}=\mathcal{O}\big{(}(1-\gamma)^{-2}\big{)}$ and $L_{h}=\mathcal{O}\big{(}(1-\gamma)^{-3}\big{)}$. Comparing with the definition of $\epsilon_{g}$ in Lemma 4.5, we choose $\eta_{0}=\frac{3\mu_{F}}{M_{g}}$ for the criteria. Thus, to satisfy $J^{\star}-\mathbb{E}[J(\theta_{K})]\leq\epsilon+\frac{\sqrt{\epsilon_{bias}}}{1-\gamma}$, we have the iteration complexity $K=\mathcal{O}({\epsilon}^{-2.5})$. In our algorithms, the sample complexity is equal to the iteration complexity, which is also $\mathcal{O}({\epsilon}^{-2.5})$.